OpenStudy (anonymous):

find the equation of the curve which satisfies the differential equation (1+y)dy+(1+x)dx=0 and passes through the origin.

OpenStudy (anonymous):

Separate and integrate. Move one of the terms to the other side and integrate both sides.

OpenStudy (anonymous):

after integrateing what i do that's a problem

OpenStudy (anonymous):

when you integrate, you'll get that +C at the end. Plug in x=0, y=0 to solve for C.

OpenStudy (anonymous):

after intergrate i get y+\[\frac{ y2 }{ 2 }\]and x+\[\frac{ x2 }{ 2}\] is it write or worng

OpenStudy (anonymous):

That's fine as long as that is all on one side of the equation. There is still the +C, so you can leave that on the other side of the equation, but you'll soon see that when you put in (0,0), C=0.

OpenStudy (anonymous):

but ans. is x^2+y^2+2x+2y=0

OpenStudy (anonymous):

Yeah, that'll work; after you have it set =0, you can multiply everything by 2 to clear the fractions.

OpenStudy (anonymous):

i can't understand can you tell me in detail

OpenStudy (anonymous):

I suppose. I'll show you my steps: \[(1+y)dy+(1+x)dx=0\] \[\int\limits (1+y)dy+ \int\limits (1+x)dx=0\] \[y+\frac{y^2}{2}+x+\frac{x^2}{2}+C=0\] \[x=0,y=0 \rightarrow 0+\frac{0^2}{2}+0+\frac{0^2}{2}+C=0 \rightarrow C=0\]

OpenStudy (anonymous):

\[\rightarrow y+\frac{y^2}{2}+x+\frac{x^2}{2}=0\] Multiply everything by 2, and the equation is: \[2y+y^2+2x+x^2=0\]

OpenStudy (anonymous):

You can verify by using implicit differentiation to find dy/dx of that equation, and you can also find dy/dx by rearranging the original differential equation, and you'll see that the two are equal.

OpenStudy (anonymous):

thanks for help me

OpenStudy (anonymous):

My pleasure.